OpenStudy (anonymous):

Find the value of x in each of the following equations. @ganeshie8

OpenStudy (anonymous):

\[6^{2x+1}=5^{x}\]

OpenStudy (anonymous):

\[2x+1=\log _{6}5^x\]

OpenStudy (alexandervonhumboldt2):

yes that corrrect

OpenStudy (anonymous):

Next step what should we do?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if you actually want an answer, you cannot do it this way you have to rewrite it in terms of either log base ten, or the natural log

OpenStudy (misty1212):

\[6^{2x+1}=5^x\\ (2x+1)\ln(x)=x\ln(5)\] then algebra

OpenStudy (misty1212):

oops typo

OpenStudy (misty1212):

\[6^{2x+1}=5^x\\ (2x+1)\ln(6)=x\ln(5)\]

OpenStudy (misty1212):

damn typos!!

OpenStudy (misty1212):

third time is a charm

OpenStudy (misty1212):

solve that equation for \(x\) using algebra remember that \(\ln(6)\) and \(\ln(5)\) are just numbers

OpenStudy (anonymous):

can u show me the working? @misty1212

OpenStudy (misty1212):

yeah it is all algebra

OpenStudy (misty1212):

\[(2x+1)\ln(6)=x\ln(5)\] distribute \[2\ln(x)x+\ln(6)=x\ln(5)\]

OpenStudy (misty1212):

subtract \(2\ln(6)\ln(x)\) from both sides \[\ln(6)=x\ln(5)-2\ln(6)x\]

OpenStudy (misty1212):

factor out the \(x\) \[\ln(6)=x(\ln(5)-2\ln(6))\]

OpenStudy (misty1212):

then divide to solve for \(x\)

OpenStudy (anonymous):

Thnx @misty1212

OpenStudy (anonymous):

Answer will be x=-0.9076

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]